OpenStudy (anonymous):

Help! Solution problems ): Difficult for me -3x - 2y = -12 9x + 6y = -9 and x - y = 11 -x + y = -11

OpenStudy (anonymous):

You can solve this by method of substitution. You do this by taking one of the equations and "isolating" either x or y. we'll start with your top problem. If you isolate x you get: x=4-(2/3)y After you isolate one of the variables you simply "substiute" what you got into the second equation. So you plug x into the second equation. If you do this you get: 9(4-(2/3)y)+6y=-9 At this point you just solve for y. When we try to simplify the left side, though, y ends up disappearing. So you get something that looks like this: 36=-9 Since we know this is not true, this problem has NO SOLUTION Doing a similar approach to your 2nd problems yields: -11=-11 This answer is true but the x and y variables disappeared! This means that the answer is "infinitely many solutions"

OpenStudy (texaschic101):

-3x - 2y = -12-->(3) 9x + 6y = -9 --------------- -9x - 6y = - 36 (result of multiplying by 3) 9x + 6y = -9 ----------------add 0 = - 45 (incorrect) I do not believe there is a solution to this problem ============================== x - y = 11 -x + y = -11 --------------add 0 + 0 = 0 This has infinite solutions